# ”Living Within Four Walls”: Exploring Emotional and Social Dynamics in Mobile Usage During Home Confinement

### v3

@@ 3.2. Procedures
Prior to data collection, participants were instructed to install the Balance for Android and Balance for Desktop apps(27; 80) on their smartphones and desktops. These apps, designed to record usage behaviors and enable experience sampling, were optimized for minimal computational, energy, and storage consumption. The apps used Experience Sampling Method (ESM) to prompt participants with a short questionnaire every 90 minutes during active device use between 7 am and 10 pm. To minimize participant burden, surveys were only triggered during active use. Additionally, the event-based ESM was triggered after 10 minutes of interaction, with follow-ups every 30 minutes. At day’s end, participants received an end-of-day survey via email at 7:30 pm. After the data collection, each participant was compensated with a 25 AUD gift card as a token of appreciation for their time.
Privacy, ethics, and safety. Our field study was designed and conducted after carefully reviewing potential risks and benefits, in compliance with the city’s COVID-19 safety regulations. The study has been approved by the Human Research Ethics Committee at the authors’ university. To ensure privacy, all personal identifying information was anonymized before data analysis. Participants’ consent was obtained prior to data collection, with a clear explanation of the study’s purpose and the nature of the data being collected. We also implemented strong data security measures, including encryption of sensitive data, to further protect participants’ privacy. Participants were informed of their right to withdraw at any time without penalty, and their data would be excluded from the study upon request.
Table 1. Overview of the collected data records across devices
@@
Table 2. Detailed overview of data categories, items, and response options
@@ 3.3. Collected Data
ESM responses. There are some popular questionnaires to measure well-being (34; 9), and the ESM questionnaire adapted from the Self-Assessment Manikin (SAM) (9) was employed in this research. In the ESM questionnaire, we collected information on valence, arousal, social role, interruptibility, and tasks from users (see Table 2). Specifically, the measurement of items valence and arousal were adapted from SAM  [^4] Compared to the original SAM questionnaire, we added a time constraint, asking participants to describe their feelings over the past hour. and rated with a 5-point Likert scale, from unhappy to happy for valence and from calm to excited for arousal. These measures were used to represent the users’ emotions based on the circumplex model (67). The social role, interruptibility, and task questions were sourced from Anderson’s assessment framework (4).
@@
Figure 2. The screenshots for the ESM and End-of-day survey
@@
4. Understanding Emotions, Social Roles, and Work-life-balance During COVID-19 Home Confinement (RQ1)
In this section, we address RQ1 ”How are emotions and social roles during COVID-19 home confinement perceived through self-reports and inferred from app usage patterns?”. First, we analyse 1522 ESM responses collected via smartphones from 25 participants. Next, we explore the impact of the COVID-19 lockdown by examining the end-of-day survey. Additionally, we explore users’ app usage behaviours through unsupervised activity recognition, identifying common activity clusters and their relationship to the perceived psychological metrics.
4.1. The Perceived Valence, Arousal and Social Roles During Covid-19 Home Confinement
@@
Next, we investigate the role of social roles in shaping emotional experiences during home confinement. A social role refers to the expectations, norms, and behaviours associated with particular social contexts (63). These roles usually influence users’ behaviours and the type of information they prefer to receive (4). As shown in Figure 3(b), participants predominantly engaged in private-related roles, suggesting individuals tend to prioritize their personal lives and relationships during home confinement.
We further explore how emotions vary across different social roles. Figure 3(c) shows that participants reported the highest valence and lowest arousal when engaged in private-related roles, compared to work-related or hybrid roles. This suggests that personal activities, which are often less stressful and more enjoyable, foster a greater sense of well-being. Conversely, work-related and hybrid roles were associated with higher levels of stress and pressure. Welch’s t-tests (89) confirm these findings, showing significant differences in valence between private-related and the other two roles (i.e., private vs. both, private vs. work), with p-value ¡ 0.05. The findings offer insights into how individuals can structure their daily routines to boost positive emotions during home confinement. Interventions aimed at improving well-being in such situations may benefit from considering the social roles individuals are engaging in (26).
@@
Figure 3. The distribution of perceived valence, arousal, social roles and their relationship
4.2. The Influence of the Home Confinement on Emotions, Social Roles and Work-Life Balance
@@
In terms of valence, 13 participants reported negative effects, describing feelings of anxiety, sadness, and boredom. Common complaints included missing social interactions (9 responses), the inability to go outside (10 responses), and a sense of boredom (6 responses). For instance, ”I thought I might be able to see friends, but I did not”, ”I want to go outside, to the park”, ”I’m missing my family”, ”It’s hard to say how it’s specifically changed my happiness today because it’s the build-up of it has gone on for so long. I’m bored and sick of my housemates and it feels like time has stopped”. On the other hand, two participants reported positive effects, such as feeling happier or using the extra time for physical activity and seamless work transitions.
@@
Figure 4. The distribution of productivity, interruption, and influence during the lockdown
Regarding arousal, 11 participants indicated a negative impact, expressing feelings of fatigue (11 responses), lack of motivation (3 responses), or laziness (5 responses), e.g., ”I don’t feel motivated to do productive things”, ”Emotionally tired, perhaps as a result of the lockdown”. However, two participants noted that they felt more energized during the lockdown.
In terms of work-life balance, the majority of participants reported they faced increased interruptions or distractions (14 responses), they had difficulty separating work from personal life (11 responses), and they spent more time working due to these disruptions (15 responses), with comments such as ”I worked a bit more than I should’ve”, ”Everything blends together, separating work and life is really difficult”. These findings suggest that the lockdown primarily had a negative impact on participants’ work-life balance, with many struggling to manage personal and professional responsibilities. While some respondents reported positive effects, such as spending more time with family or gardening, the overall trend indicates that the lockdown negatively influenced emotional well-being and work-life balance.
4.3. Understanding App Usage Behaviours During COVID-19 Home Confinement
Table 3. App categories (referred to the Google Play Store categories) and top three frequently used apps.
@@ 4.3.1. App Categorization
4.3.2. Unsupervised Activity Recognition From App Sequences
@@
Segmentation. To investigate the relationship between app usage and various factors, we propose App-based Information Gain Temporal Segmentation (AIGTS), a method that focuses on identifying different activities within app usage. Figure 5 illustrates how activities are extracted from app usage data. AIGTS is inspired by Information Gain Temporal Segmentation (IGTS) (68), an unsupervised method for segmenting multivariate time series. IGTS identifies splits that minimize the weighted entropy across the multivariate time series, where each variable corresponds to whether a specific app category was used.
@@
Figure 5. An illustration of unsupervised activity recognition using app usage data. The Segment Vector indicates the app subcategory.
Clustering. Next, AIGTS is applied to the all-time series, resulting in a segmented list for each user. To extract features from these segments, K-means clustering (41) was applied. The distribution of different categories appearing in a segment was treated as an $n$ dimensional vector, where $n$ is the number of app categories. To identify the optimal number of segments, Silhouette Analysis (42) was utilized with a maximum of 20 clusters. Especially, the label indicating which cluster a segment belongs to was used as a feature representing the activity within that segment. These segment labels were derived manually from the average distribution of apps in that cluster. In this paper, a segment that belongs to a particular cluster will be referred to as an ‘activity’.
Figure 6. Proportion of app subcategory usage across user activity clusters. Clusters were manually named to describe the distribution of average activities within each cluster.
@@
The Device Idle cluster, corresponding to Personalization apps (e.g., launchers, as shown in Table 3), represents periods when users remain on their device’s home screen. We also found that the clusters Communication Only and Device Idle were the most frequent user activities during the COVID-19 home confinement, accounting for 17% and 14.6% of all activities, respectively. These activities are characterized by high proportions of communication-related apps and personalization apps, suggesting users spent considerable time on their devices in passive states or communicating with others. In addition, clusters such as Work-Related Communication and Hybrid Work exhibited higher proportions of productivity and business apps. These clusters likely reflect professional tasks and remote work activities, which were common during the lockdown period. The differentiation between Light Work and Deep Work clusters highlights variations in the intensity and type of work-related app usage.
4.3.3. Correlation Analysis of User Activity/App Usage and Well-Being
. We further investigate how user activities and app usage behaviours correlate with four key dimensions of well-being: valence, arousal, productivity, and interruptions. To assess these correlations, we performed Spearman Rank Correlation (18) tests between the time spent on different activities/apps and the scores for each dimension. Specifically, valence and arousal were examined for theirr relationship to the time spent on user activities in the preceding hour, while productivity and interruptions were calculated based on the end-of-day surveys.
Table 4. Top correlations between app usage/user acitivty and well-being dimensions.
@@ 4.3.3. Correlation Analysis of User Activity/App Usage and Well-Being
Interestingly, as shown in Table 4, we found mostly weak correlations, with a few activities exhibiting moderate relationships. For instance, activities related to Tools Apps (e.g., phone home screens and keyboards) showed a moderate positive correlation with valence, especially among older adults. This suggests that certain habitual or passive tasks, such as interacting with productivity tools, may positively influence mood, possibly linked to work-related or routine activities. Additionally, Surfing activity (e.g., browsing news, including COVID-19 updates) showed a negative correlation with valence among women and the younger age group, indicating that news consumption during this period may be linked to lower mood. The correlation with productivity was notably negative for Communication Apps (e.g., keyboards, chat apps), but positively correlated with Work-Related Communication activity (e.g., business apps like Microsoft Teams), highlighting that whether a communication tool is used for professional or personal purposes may influence productivity levels.
@@
5. Predicting User Emotion and Social Roles from Mobile Usage Data (RQ2)
In this section, we address RQ2 ”Can mobile usage data be used to model emotional well-being during home confinement, and which features are most predictive?”. First, we introduce the features extracted from mobile usage data and the prediction pipeline. We then list and discuss the prediction results of the general models and individual models. Finally, we examine the significant features for predicting human well-being.
Table 5. The computed features from smartphone usage behaviours. For the valence model, the features selected by F-statistic $\geq 5^{*}$ or $\geq 10^{**}$. For the arousal model, the features selected by F-statistic $\geq 5^{+}$ or $\geq 10^{++}$.
@@ 5. Predicting User Emotion and Social Roles from Mobile Usage Data (RQ2)
5.1. Feature Extraction
Table 5 provides an overview of the features utilized in our research. Similar to previous mobile sensing studies (40; 69; 86), we extract contextual information, notifications-related information, physical activity, app, and activities inferred from apps. To process the app-usage data, we divide it into segments based on the ESM recording frequency (set at 90 minutes) and stored it in temporal sequences. Subsequently, we extract the features from these 90-minute segments. In order to handle cyclical features such as days (e.g., Monday) and times (e.g., Morning), we transform them into sine and cosine values. Additionally, it is worth noting that app usage patterns exhibit variations between weekdays and weekends (35; 40), prompting us to conduct separate analyses for these two categories. Regarding the time period per day, we divide it into six 4-hour segments, i.e., Early Morning, Morning, Afternoon, Evening, Night, and Late Night. Certain frequency-based features, such as notification-related or physical-activity-related features, are converted using quartiles, specifically the minimum, 25th percentile, 50th percentile, and 75th percentile. For instance, the number of interacted notifications is categorized into groups as ‘0’, ‘1-3’, ‘4-9’, and ‘10+’. Furthermore, the physical activities considered in our study only include active physical activities like bicycling, walking, and running.
To capture user activities related to app usage, we employed our proposed method, AIGTS, as detailed in Section 4.3.2. AIGTS allows us to extract both activities and their corresponding durations. Based on the extracted information, we compute the feature ‘entropy_user_activity’ by determining the proportion of each activity’s time portion within the entire 90-minute time interval. In addition to extracting individual features such as ‘num_noti’, we also generate composite features by combining two existing features. For example, the feature ‘num_communication’ is calculated by selecting and combining the notification sources. This composite feature counts the number of notifications originating the Business, Communication, or Social app categories. Notifications that do not belong to these app categories are counted as ‘app_noti’.
5.2. Prediction Pipeline
While emotion prediction can be addressed as a classification problem, with emotion levels categorized into two or three classes based on predefined thresholds (58; 88). We believe that utilizing regression is a better approach because valence and arousal labels are inherently ordinal data, where the relative order holds significance. In this research, we establish a regression-based pipeline to forecast the emotion scores of participants, as detailed below.
@@ 5.2. Prediction Pipeline
In this project, we aim to develop regression models that accurately predict human emotions during home confinement. To accomplish this, we employ three commonly used regression models: K-Nearest Neighbor (KNN) (72), Random Forest (RF) (71), and Gradient Boosting (GB) (21). KNN is a widely utilized model known for its versatility across various domains. It works by finding the K nearest neighbours to a given data point and using their values to predict the target variable. RF and GB, on the other hand, are ensemble models based on decision trees. These models combine predictions from multiple trees. However, there is a distinction in their approach. GB constructs trees or aggregates results in a way that minimizes the loss function, whereas RF constructs individual trees or aggregates results simultaneously.
@@
Various methods, such as SHAP (44), LIME (65), and the $F$-statistic ($F$) (60), can be used for feature selection. In this study, we employ the $F$-statistic to determine the appropriate number of features. While most of our features are numeric, the $F$-statistic measures the univariate linear dependence between each feature and the target variable, taking into account statistical significance. Generally, a higher $F$ value indicates greater importance of the feature in predicting the target variable. In our experiments, we compare regression models using three feature sets: all features, features selected based on an $F$ value greater than or equal to 5, and greater than or equal to 10.
@@
To address the limited richness of our data, we implemented Nested Cross-Validation (CV) (50) to train and evaluate our prediction models. Nested CV comprises an outer and an inner loop, both employing k-fold cross-validation (83) with different values of $k$. The outer loop splits the data into $k$ folds, using each fold for testing. The inner loop further divides the remaining data into $k$ folds, with each fold serving as a validation set to train the model and optimize hyperparameters. The best parameter settings are selected based on the inner loop results. This approach provides an accurate estimate of the model’s unbiased generalization performance.
@@
Baselines and Metrics. Similar to previous human-centered studies (24; 82), we compare the proposed models with two baselines, Mean, which calculates the average score of user responses, and Median, which calculates the median value based on the data distribution and treats it as a predicted value. To evaluate the performance of regression models, we employ two commonly used metrics: Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) (13). Generally, lower values of MAE or RMSE indicate better prediction performance.
5.3. Prediction Result
5.3.1. General Model

### v4

@@ 3.2 Procedures
Prior to data collection, participants were instructed to install the Balance for Android and Balance for Desktop apps[60, 61] on their smartphones and desktops. These apps, designed to record usage behaviors and enable experience sampling, were optimized for minimal computational, energy, and storage consumption. The apps used Experience Sampling Method (ESM) to prompt participants with a short questionnaire every 90 minutes during active device use between 7 am and 10 pm. To minimize participant burden, surveys were only triggered during active use. Additionally, the event-based ESM was triggered after 10 minutes of interaction, with follow-ups every 30 minutes. At day’s end, participants received an end-of-day survey via email at 7:30 pm. After the data collection, each participant was compensated with a 25 AUD gift card as a token of appreciation for their time.
Privacy, ethics, and safety.
Our field study was designed and conducted after carefully reviewing potential risks and benefits, in compliance with the city’s COVID-19 safety regulations. The study has been approved by the Human Research Ethics Committee at the authors’ university. To ensure privacy, all personal identifying information was anonymized before data analysis. Participants’ consent was obtained prior to data collection, with a clear explanation of the study’s purpose and the nature of the data being collected. We also implemented strong data security measures, including encryption of sensitive data, to further protect participants’ privacy. Participants were informed of their right to withdraw at any time without penalty, and their data would be excluded from the study upon request.
Table 1: Overview of the collected data records across devices
@@
Table 2: Detailed overview of data categories, items, and response options
@@ 3.3 Collected Data
ESM responses. There are some popular questionnaires to measure well-being [62, 63], and the ESM questionnaire adapted from the Self-Assessment Manikin (SAM) [63] was employed in this research. In the ESM questionnaire, we collected information on valence, arousal, social role, interruptibility, and tasks from users (see Table 2). Specifically, the measurement of items valence and arousal were adapted from SAM  [^7] Compared to the original SAM questionnaire, we added a time constraint, asking participants to describe their feelings over the past hour. and rated with a 5-point Likert scale, from unhappy to happy for valence and from calm to excited for arousal. These measures were used to represent the users’ emotions based on the circumplex model [16]. The social role, interruptibility, and task questions were sourced from Anderson’s assessment framework [64].
@@
Figure 2: The screenshots for the ESM and End-of-day survey
@@
4 Understanding Emotions, Social Roles, and Work-life-balance During Home Confinement (RQ1)
In this section, we address RQ1 “How are emotions and social roles during home confinement perceived through self-reports and inferred from app usage patterns?”. First, we analyse 1522 ESM responses collected via smartphones from 25 participants. Next, we explore the impact of the lockdown by examining the end-of-day survey. Additionally, we explore users’ app usage behaviours through unsupervised activity recognition, identifying common activity clusters and their relationship to the perceived psychological metrics.
4.1 The Perceived Valence, Arousal and Social Roles During Home Confinement
@@
Figure 3: The distribution of perceived valence, arousal, social roles and their relationship
Next, we investigate the role of social roles in shaping emotional experiences during home confinement. A social role refers to the expectations, norms, and behaviours associated with particular social contexts [65]. These roles usually influence users’ behaviours and the type of information they prefer to receive [64]. As shown in Figure 3(b), participants predominantly engaged in private-related roles, suggesting individuals tend to prioritize their personal lives and relationships during home confinement.
[…]
4.2 The Influence of the Home Confinement on Emotions, Social Roles and Work-Life Balance
@@
In terms of valence, 13 participants reported negative effects, describing feelings of anxiety, sadness, and boredom. Common complaints included missing social interactions (9 responses), the inability to go outside (10 responses), and a sense of boredom (6 responses). For instance, “I thought I might be able to see friends, but I did not”, “I want to go outside, to the park”, “I’m missing my family”, “It’s hard to say how it’s specifically changed my happiness today because it’s the build-up of it has gone on for so long. I’m bored and sick of my housemates and it feels like time has stopped”. On the other hand, two participants reported positive effects, such as feeling happier or using the extra time for physical activity and seamless work transitions.
Regarding arousal, 11 participants indicated a negative impact, expressing feelings of fatigue (11 responses), lack of motivation (3 responses), or laziness (5 responses), e.g., “I don’t feel motivated to do productive things”, “Emotionally tired, perhaps as a result of the lockdown”. However, two participants noted that they felt more energized during the lockdown.
In terms of work-life balance, the majority of participants reported they faced increased interruptions or distractions (14 responses), they had difficulty separating work from personal life (11 responses), and they spent more time working due to these disruptions (15 responses), with comments such as “I worked a bit more than I should’ve”, “Everything blends together, separating work and life is really difficult”. These findings suggest that the lockdown primarily had a negative impact on participants’ work-life balance, with many struggling to manage personal and professional responsibilities. While some respondents reported positive effects, such as spending more time with family or gardening, the overall trend indicates that the lockdown negatively influenced emotional well-being and work-life balance.
@@
Figure 4: The distribution of productivity, interruption, and influence during the lockdown
4.3 Understanding App Usage Behaviours During Home Confinement
Table 3: App categories (referred to the Google Play Store categories) and top three frequently used apps.
@@ 4.3.1 App Categorization
4.3.2 Unsupervised Activity Recognition From App Sequences
@@
Segmentation. To investigate the relationship between app usage and various factors, we propose App-based Information Gain Temporal Segmentation (AIGTS), a method that focuses on identifying different activities within app usage. Figure 5 illustrates how activities are extracted from app usage data. AIGTS is inspired by Information Gain Temporal Segmentation (IGTS) [68], an unsupervised method for segmenting multivariate time series. IGTS identifies splits that minimize the weighted entropy across the multivariate time series, where each variable corresponds to whether a specific app category was used.
@@
Figure 5: An illustration of unsupervised activity recognition using app usage data. The Segment Vector indicates the app subcategory.
Clustering. Next, AIGTS is applied to the all-time series, resulting in a segmented list for each user. To extract features from these segments, K-means clustering [69] was applied. The distribution of different categories appearing in a segment was treated as an $n$ dimensional vector, where $n$ is the number of app categories. To identify the optimal number of segments, Silhouette Analysis [70] was utilized with a maximum of 20 clusters. Especially, the label indicating which cluster a segment belongs to was used as a feature representing the activity within that segment. These segment labels were derived manually from the average distribution of apps in that cluster. In this paper, a segment that belongs to a particular cluster will be referred to as an ‘activity’.
Figure 6: Proportion of app subcategory usage across user activity clusters. Clusters were manually named to describe the distribution of average activities within each cluster.
@@
The Device Idle cluster, corresponding to Personalization apps (e.g., launchers, as shown in Table 3), represents periods when users remain on their device’s home screen. We also found that the clusters Communication Only and Device Idle were the most frequent user activities during the home confinement, accounting for 17% and 14.6% of all activities, respectively. These activities are characterized by high proportions of communication-related apps and personalization apps, suggesting users spent considerable time on their devices in passive states or communicating with others. In addition, clusters such as Work-Related Communication and Hybrid Work exhibited higher proportions of productivity and business apps. These clusters likely reflect professional tasks and remote work activities, which were common during the lockdown period. The differentiation between Light Work and Deep Work clusters highlights variations in the intensity and type of work-related app usage.
4.3.3 Correlation Analysis of User Activity/App Usage and Well-Being
. We further investigate how user activities and app usage behaviors correlate with four key dimensions of well-being: valence, arousal, productivity, and interruptions. To assess these correlations, we performed Spearman Rank Correlation [71] tests between the time spent on different activities/apps and the scores for each dimension. Specifically, valence and arousal were examined for their relationship to the time spent on user activities in the preceding hour, while productivity and interruptions were calculated based on the end-of-day surveys.
Table 4: Top correlations between app usage/user activity and well-being dimensions.
@@ 4.3.3 Correlation Analysis of User Activity/App Usage and Well-Being
Interestingly, as shown in Table 4, we found mostly weak correlations, with a few activities exhibiting moderate relationships. For instance, activities related to Tools Apps (e.g., phone home screens and keyboards) showed a moderate positive correlation with valence, especially among older adults. This suggests that certain habitual or passive tasks, such as interacting with productivity tools, may positively influence mood, possibly linked to work-related or routine activities.
Additionally, Surfing activity (e.g., browsing news, including pandemic updates) showed a negative correlation with valence among women and the younger age group, indicating that news consumption during this period may be linked to lower mood. The correlation with productivity was notably negative for Communication Apps (e.g., keyboards, chat apps), but positively correlated with Work-Related Communication activity (e.g., business apps like Microsoft Teams), highlighting that whether a communication tool is used for professional or personal purposes may influence productivity levels.
@@
5 Predicting User Emotion and Social Roles from Mobile Usage Data (RQ2)
In this section, we address RQ2 “Can mobile usage data be used to model emotional well-being during home confinement, and which features are most predictive?”. First, we introduce the features extracted from mobile usage data and the prediction pipeline. We then list and discuss the prediction results of the general models and individual models. Finally, we examine the significant features for predicting human well-being.
Table 5: The computed features from smartphone usage behaviours. For the valence model, the features selected by F-statistic $\geq 5^{*}$ or $\geq 10^{**}$. For the arousal model, the features selected by F-statistic $\geq 5^{+}$ or $\geq 10^{++}$.
@@ 5 Predicting User Emotion and Social Roles from Mobile Usage Data (RQ2)
5.1 Feature Extraction
Table 5 provides an overview of the features utilized in our research. Similar to previous mobile sensing studies [72, 73, 74], we extract contextual information, notifications-related information, physical activity, app, and activities inferred from apps. To process the app-usage data, we divide it into segments based on the ESM recording frequency (set at 90 minutes) and store it in temporal sequences. Subsequently, we extract the features from these 90-minute segments. In order to handle cyclical features such as days (e.g., Monday) and times (e.g., Morning), we transform them into sine and cosine values. Additionally, it is worth noting that app usage patterns exhibit variations between weekdays and weekends [75, 72], prompting us to conduct separate analyses for these two categories. Regarding the time period per day, we divide it into six 4-hour segments, i.e., Early Morning, Morning, Afternoon, Evening, Night, and Late Night. Certain frequency-based features, such as notification-related or physical-activity-related features, are converted using quartiles, specifically the minimum, 25th percentile, 50th percentile, and 75th percentile. For instance, the number of interacted notifications is categorized into groups as ‘0’, ‘1-3’, ‘4-9’, and ‘10+’. Furthermore, the physical activities considered in our study only include active physical activities like bicycling, walking, and running.
To capture user activities related to app usage, we employed our proposed method, AIGTS, as detailed in Section 4.3.2. AIGTS allows us to extract both activities and their corresponding durations. Based on the extracted information, we compute the feature ‘entropy_user_activity’ by determining the proportion of each activity’s time portion within the entire 90-minute time interval. In addition to extracting individual features such as ‘num_noti’, we also generate composite features by combining two existing features. For example, the feature ‘num_communication’ is calculated by selecting and combining the notification sources. This composite feature counts the number of notifications originating from the Business, Communication, or Social app categories. Notifications that do not belong to these app categories are counted as ‘app_noti’.
5.2 Prediction Pipeline
While emotion prediction can be addressed as a classification problem, with emotion levels categorized into two or three classes based on predefined thresholds [76, 42]. We believe that utilizing regression is a better approach because valence and arousal labels are inherently ordinal data, where the relative order holds significance. In this research, we establish a regression-based pipeline to forecast the emotion scores of participants, as detailed below.
@@ 5.2 Prediction Pipeline
In this project, we aim to develop regression models that accurately predict human emotions during home confinement. To accomplish this, we employ three commonly used regression models: K-Nearest Neighbor (KNN) [77], Random Forest (RF) [78], and Gradient Boosting (GB) [79]. KNN is a widely utilized model known for its versatility across various domains. It works by finding the K nearest neighbors to a given data point and using their values to predict the target variable. RF and GB, on the other hand, are ensemble models based on decision trees. These models combine predictions from multiple trees. However, there is a distinction in their approach. GB constructs trees or aggregates results in a way that minimizes the loss function, whereas RF constructs individual trees or aggregates results simultaneously.
@@
Various methods, such as SHAP [80], LIME [81], and the $F$-statistic ($F$) [82], can be used for feature selection. In this study, we employ the $F$-statistic to determine the appropriate number of features. While most of our features are numeric, the $F$-statistic measures the univariate linear dependence between each feature and the target variable, taking into account statistical significance. Generally, a higher $F$ value indicates greater importance of the feature in predicting the target variable. In our experiments, we compare regression models using three feature sets: all features, features selected based on an $F$ value greater than or equal to 5, and greater than or equal to 10.
@@
To address the limited richness of our data, we implemented Nested Cross-Validation (CV) [83] to train and evaluate our prediction models. Nested CV comprises an outer and an inner loop, both employing k-fold cross-validation [84] with different values of $k$. The outer loop splits the data into $k$ folds, using each fold for testing. The inner loop further divides the remaining data into $k$ folds, with each fold serving as a validation set to train the model and optimize hyperparameters. The best parameter settings are selected based on the inner loop results. This approach provides an accurate estimate of the model’s unbiased generalization performance.
@@
Baselines and Metrics. Similar to previous human-centered studies [44, 85], we compare the proposed models with two baselines, Mean, which calculates the average score of user responses, and Median, which calculates the median value based on the data distribution and treats it as a predicted value. To evaluate the performance of regression models, we employ two commonly used metrics: Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) [86]. Generally, lower values of MAE or RMSE indicate better prediction performance.
5.3 Prediction Result
5.3.1 General Model